ax-2y=4=0 we know that f(-2)=-8 what is the value of a?
Answer by xcentaur(357)

ax-2y=4=0
How can the same expression equal 4 and 0?
Is the question correct,and have you typed it out exactly as it is?


Still,I will guess here:
ax-2y+4=0
f(-2)=-8
a(-2)-2y+4=8
-2a-2y+4-8=0
-2a-2y-4=0
-2a=4+2y
a=-2-y


If this answer does not match with the provided solution,please recheck your question and post again